# A Continuous-Time Ensemble Kalman–Bucy Smoother for Causal Inference and Model Discovery

### 28 Apr 2026 v1

@@ Abstract
keywords:
Continuous-time data assimilation , Ensemble Kalman–Bucy filter and smoother , Parameter estimation , Causal inference , Model identification
@@ 1 Introduction
Complex dynamical systems arise in many scientific areas, such as atmospheric and ocean science, fluid dynamics, geophysics, and neuroscience [1, 2, 3, 4, 5]. Such systems are characterized by high dimensionality, strong nonlinearity, and nontrivial driving mechanisms [3, 6, 7, 8].
@@
A central mathematical problem when studying complex dynamical systems is reconstructing the full state of the system from partial and noisy observations, which is essential for system reanalysis, uncertainty quantification, forecasting, and optimal control [9, 10, 11, 12, 13]. More importantly, accurate state estimation is often an essential prerequisite for model identification and the discovery of the underlying causal structure [14, 15, 16].
@@
However, accurate state estimation is challenging. Uncertainties arise primarily from two sources: measurement noise due to imperfect observation tools or methodologies, and model error stemming from incomplete physical understanding or limited computational resolution. In chaotic systems, these uncertainties are amplified over time through multiscale nonlinear interactions, so state estimates that rely entirely on model simulations can quickly become unreliable [11, 17, 18]. In contrast, purely data-driven approaches that rely on statistical interpolation of time-series data often fail to incorporate temporal dynamical knowledge and to enforce physics constraints [15, 19]. A strategy is therefore needed that integrates observational information with model dynamics while addressing the limitations of each. Data assimilation (DA) has been the de facto approach, leading to a powerful framework that systematically integrates observations with dynamics to improve state estimation and reduce uncertainty and biases [20, 21, 12]. Over the past few decades, DA and its variants have been extensively developed and deployed across various disciplines, with notable successes in numerical weather prediction and medical imaging [22, 23, 24].
@@
Broadly speaking, classical DA algorithms are categorized into filters and smoothers. Filters estimate the state using observations up to and including the present, whereas smoothers incorporate information from the entire observation window, including past, present, and future observations. While filtering is the standard approach for real-time forecasting, smoothing is theoretically more accurate for retrospective analysis [25, 26, 27]. On the one hand, smoothing systematically improves estimation error and reduces uncertainty [25, 28]. On the other hand, smoothing can mitigate the detection delay that filters often exhibit at the onset of extreme events. This distinction becomes particularly important in systems where the underlying dynamics are driven by hidden forcing mechanisms. As a result, filters tend to detect the driving signal only after its effects have been observed [29]. In contrast, by leveraging future information, smoothers can more accurately identify the onset and persistence of the underlying driving mechanisms [16, 30, 31].
@@
Although a wide range of DA algorithms has been developed, many methods face theoretical and practical limitations. The Kalman filter and the Rauch–Tung–Striebel (RTS) smoother provide optimal solutions, but only for linear-Gaussian systems [32, 33]. In complex dynamical systems, however, strongly non-Gaussian features such as skewness and heavy tails often violate the Gaussian assumptions underlying the classical theory [34]. For general nonlinear dynamics, variational DA approaches typically provide deterministic state estimates that do not quantify uncertainty, while particle methods suffer from severe computational burdens in high dimensions due to the curse of dimensionality [35, 36, 37]. Consequently, ensemble Kalman methods have emerged as an effective alternative [38, 39]. By approximating system statistics with a finite ensemble, these methods provide state estimates that explicitly quantify uncertainty while remaining derivative-free and computationally feasible in many high-dimensional settings. In practice, the ensemble size (typically $O(10^{2}))$ is often chosen independently of the system’s effective dimension [40, 41]. Due to these advantages, ensemble methods have been widely adopted in large-scale operational applications, including numerical weather prediction, ocean science, geophysics, and fluid dynamics [42, 43, 40, 44]. Most existing ensemble filtering and smoothing algorithms, however, are formulated in discrete time, reflecting the sequential structure of forecast-analysis cycles. Efficient continuous-time smoothers for nonlinear systems remain comparatively less explored, despite their importance for systems naturally described by stochastic differential equations. Existing formulations, such as the continuous-time unscented RTS smoother [45], directly propagate the smoother mean and covariance. In contrast, our ensemble-based smoother evolves stochastic trajectories, providing conditional samples for downstream tasks such as model discovery, while naturally admitting covariance localization for high-dimensional systems.
@@
In this paper, we develop a continuous-time ensemble Kalman–Bucy smoother (EnKBS) for efficient state estimation and trajectory sampling in complex nonlinear dynamical systems. To address high-dimensional settings, the method incorporates standard regularization techniques, including covariance localization and inflation, which mitigate sampling errors, improve numerical stability, and prevent ensemble divergence in nonlinear models [46, 47, 48]. Beyond state estimation, we demonstrate that EnKBS provides a powerful tool for dynamical discovery. We incorporate EnKBS into a Bayesian causal inference framework that identifies causal relationships by tracing driving signals backwards from their observed effects [16]. Numerical tests are presented using a nonlinear dyad model exhibiting extreme events, for which closed-form filtering and smoothing distributions are available [16, 49]. We verify asymptotic convergence of the method to the exact analytical solution and show that it remains accurate and statistically consistent even with an ensemble size of $O(10)$. We further integrate EnKBS into a causality-based iterative learning algorithm for partially observed systems, where EnKBS-based trajectory sampling enables the recovery of hidden parameters and identification of the underlying model structure [15].
@@
The rest of this paper is organized as follows. Section 2 formulates the problem and introduces the proposed continuous-time EnKBS, together with regularization strategies for high-dimensional systems. Section 3 evaluates EnKBS on the high-dimensional Lorenz-96 system [18] and presents applications to assimilative causal inference (ACI) and causality-based model discovery [15, 16]. We conclude the paper with a discussion in Section 4. Mathematical derivations of the EnKBS and its connection to the discrete-time RTS smoother in a formal time limit are deferred to A. Finally, B shows that the proposed EnKBS recovers the exact mean and covariance equations of the classical continuous-time RTS smoother in the linear-Gaussian setting.
@@ 2.1 Problem Setup
We consider a general class of continuous-time nonlinear systems. Let $\mathbf{x}(t)\in\mathbb{R}^{n_{\mathbf{x}}}$ denote the hidden states and $\mathbf{y}(t)\in\mathbb{R}^{n_{\mathbf{y}}}$ denote the observations (data). The dynamics are governed by the following stochastic differential equations (SDEs) [34, 50]:
$\,\mathrm{d}\mathbf{x}$ $=\mathbf{f}(\mathbf{x},\mathbf{y},t)\,\mathrm{d}t+\mathbf{\Sigma}(\mathbf{y},t)^{1/2}\,\mathrm{d}\mathbf{B}(t),$ (1a)
$\,\mathrm{d}\mathbf{y}$ $=\mathbf{h}(\mathbf{x},\mathbf{y},t)\,\mathrm{d}t+\mathbf{\Gamma}(\mathbf{y},t)^{1/2}\,\mathrm{d}\mathbf{W}(t).$ (1b)
$\mathbf{f}$ is the drift of the signal process (1a) and $\mathbf{h}$ is the drift of the observation process (1b). The terms $\mathbf{B}(t)\in\mathbb{R}^{n_{\mathbf{x}}}$ and $\mathbf{W}(t)\in\mathbb{R}^{n_{\mathbf{y}}}$ are independent standard Wiener processes. For simplicity, we assume that the signal process noise covariance $\mathbf{\Sigma}(\mathbf{y},t)\in\mathbb{R}^{n_{\mathbf{x}}\times n_{\mathbf{x}}}$ and the observation noise covariance $\mathbf{\Gamma}(\mathbf{y},t)\in\mathbb{R}^{n_{\mathbf{y}}\times n_{\mathbf{y}}}$ are symmetric and positive definite. Throughout the paper, all SDEs, whether written in differential form or in formal derivative form, are understood in the Itô sense [34, 51].
We consider a fixed time horizon $t\in[0,T]$, and aim to characterize the conditional distribution of $\mathbf{x}(t)$ given the available observations. Let $\mathbf{y}_{[0,\tau]}=\{\mathbf{y}(s):0\leq s\leq\tau\}$ denote the observation trajectory up to time $\tau\in[0,T]$. We formulate the filtering and smoothing problems of DA as follows:
@@
    Filtering: Estimate the conditional distribution of $\mathbf{x}(t)$ given the observational history up to the present $\mathbf{y}_{[0,t]}$. The filtering distribution is denoted by $p(\mathbf{x}(t)\mid\mathbf{y}_{[0,t]})$, with mean $\overline{\mathbf{x}}_{\mathrm{f}}(t)$ and covariance $\mathbf{P}_{\mathrm{f}}(t)$.
@@
    Smoothing: Estimate the conditional distribution of $\mathbf{x}(t)$ given the full observation trajectory $\mathbf{y}_{[0,T]}$, including the past and the future. The smoothing distribution is denoted by $p(\mathbf{x}(t)\mid\mathbf{y}_{[0,T]})$, with mean $\overline{\mathbf{x}}_{\mathrm{s}}(t)$ and covariance $\mathbf{P}_{\mathrm{s}}(t)$.
@@
It is worth mentioning that the coefficients in (1), including the drifts $\mathbf{f}$ and $\mathbf{h}$, as well as the noise covariances $\mathbf{\Sigma}$ and $\mathbf{\Gamma}$, may depend explicitly on the observation history $\mathbf{y}_{[0,t]}$. Since $\mathbf{y}$ is given in the filter and smoother, its trajectory can be treated as time-varying parameters [52]. This flexibility allows for broad application in which the observed variables actively feed back on and modulate the hidden dynamics; however, for our purposes, we assume the Markovian structure in (1) [51]. For notational simplicity, we may omit the explicit dependence of the model components on the dependent and time variables when the context is clear, and write $\mathbf{f}(\mathbf{x},\mathbf{y},t)$ as $\mathbf{f}$, and similarly for $\mathbf{h}$, $\mathbf{\Sigma}$, and $\mathbf{\Gamma}$.
@@
In general, evolving the full filtering or smoothing distribution is computationally infeasible [52]. Therefore, we focus on estimating the first two moments: the mean and covariance. The idea is that in systems with linear dynamics and Gaussian noise, the conditional distributions are Gaussian and fully described by their mean and covariance, i.e., $p(\mathbf{x}(t)\mid\mathbf{y}_{[0,t]})=\mathcal{N}(\overline{\mathbf{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t))$ and $p(\mathbf{x}(t)\mid\mathbf{y}_{[0,T]})=\mathcal{N}(\overline{\mathbf{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t))$. These exact moments are governed by the classical Kalman–Bucy Filter [53] and continuous-time Rauch–Tung–Striebel (RTS) smoother equations [33].
@@
For general nonlinear systems where closed-form DA solutions are unavailable, we adopt an ensemble-based approach [21, 38]. Here, the conditional distributions are represented by a finite ensemble of size $m$. This can be viewed as a Monte Carlo approximation in which the conditional moments are replaced by empirical ensemble moments. Each ensemble member independently evolves according to (1), but the ensemble constituents are coupled through the empirical mean and covariance.
@@
Let $\mathbf{x}^{(i)}(t)$, $i=1,\dots,m$, denote the ensemble members. The empirical mean $\overline{\mathbf{x}}(t)$ and covariance $\mathbf{P}(t)$ for $t\in[0,T]$ are defined by
$\overline{\mathbf{x}}(t):=\frac{1}{m}\sum_{i=1}^{m}\mathbf{x}^{(i)}(t),\qquad\mathbf{P}(t):=\frac{1}{m-1}\sum_{i=1}^{m}\big(\mathbf{x}^{(i)}(t)-\overline{\mathbf{x}}(t)\big)\big(\mathbf{x}^{(i)}(t)-\overline{\mathbf{x}}(t)\big)^{\top}.$
@@
When needed, we use subscripts $\mathrm{f}$ and $\mathrm{s}$ to distinguish between the filter- and smoother-based ensembles. For brevity, we use $\big(\overline{\mathbf{x}}_{\mathrm{f}},\mathbf{P}_{\mathrm{f}}\big)$ and $\big(\overline{\mathbf{x}}_{\mathrm{s}},\mathbf{P}_{\mathrm{s}}\big)$ to denote the corresponding empirical moments, and do not distinguish them from the true moments unless confusion may arise.
@@
While filtering is the standard approach for online state estimation, many applications, such as Earth system reanalysis [11] and parameter estimation [54, 55], necessitate a retrospective analysis once the full observation trajectory over $[0,T]$, $\mathbf{y}_{[0,T]}$, is available. Based on the filtering quantities $\big(\overline{\mathbf{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t)\big)$, the smoothing quantities $\big(\overline{\mathbf{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t)\big)$ incorporate information from future observations $\mathbf{y}_{[t,T]}$, producing more accurate estimates and reducing uncertainty over $[0,T]$. This improvement motivates a systematic forward-backward structure algorithm [21, 56]:
@@
    Forward pass (filtering): Integrate forward from $t=0$ to $t=T$ to compute the filter ensemble $\mathbf{x}^{(i)}_{\mathrm{f}}(t),i=1,\dots,m$ (forecasting). The filtering mean and covariance $\{\overline{\mathbf{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t)\}_{0\leq t\leq T}$ are obtained.
@@
    Backward pass (smoothing): Integrate backward-in-time from $t=T$ to $t=0$. Using the filtering estimates as a prior, this pass produces the smoothing ensemble $\mathbf{x}_{\mathrm{s}}^{(i)}(t),i=1,\dots,m$ and therefore its statistics $\{\overline{\mathbf{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t)\}_{0\leq t\leq T}$ (hindcasting).
In our method, we implement this architecture by utilizing the continuous-time ensemble Kalman–Bucy filter (EnKBF) for the forward pass [57], and the EnKBS for the backward pass. The operational details of these methods are presented in the following subsections.

### 03 May 2026 v2

@@ Abstract
Complex dynamical systems arise in many scientific areas, such as atmospheric and ocean science, fluid dynamics, geophysics, and neuroscience (Chen, 2025; Majda, 2016; Strogatz, 2024; Izhikevich, 2006; Holmes et al., 2012). Such systems are characterized by high dimensionality, strong nonlinearity, and nontrivial driving mechanisms (Strogatz, 2024; Andreou and Chen, 2024b; Frisch, 1995; Bar-Yam, 2019).
@@
A central mathematical problem when studying complex dynamical systems is reconstructing the full state of the system from partial and noisy observations, which is essential for system reanalysis, uncertainty quantification, forecasting, and optimal control (Chen et al., 2025; Majda and Harlim, 2012a; Kalnay, 2003; Asch et al., 2016; Stengel, 1994). More importantly, accurate state estimation is often an essential prerequisite for model identification and the discovery of the underlying causal structure (Brunton et al., 2016; Chen and Zhang, 2023; Andreou et al., 2026).
@@
However, accurate state estimation is challenging. Uncertainties arise primarily from two sources: measurement noise due to imperfect observation tools or methodologies, and model error stemming from incomplete physical understanding or limited computational resolution. In chaotic systems, these uncertainties are amplified over time through multiscale nonlinear interactions, so state estimates that rely entirely on model simulations can quickly become unreliable (Kalnay, 2003; Lorenz, 1969, 2006). In contrast, purely data-driven approaches that rely on statistical interpolation of time-series data often fail to incorporate temporal dynamical knowledge and to enforce physics constraints (Chen and Zhang, 2023; Raissi et al., 2019). A strategy is therefore needed that integrates observational information with model dynamics while addressing the limitations of each. Data assimilation (DA) has been the de facto approach, leading to a powerful framework that systematically integrates observations with dynamics to improve state estimation and reduce uncertainty and biases (Law et al., 2015; Evensen, 2009; Asch et al., 2016). Over the past few decades, DA and its variants have been extensively developed and deployed across various disciplines, with notable successes in numerical weather prediction and medical imaging (Bauer et al., 2015; Lorenc, 1986; Tarantola, 2005).
@@
Broadly speaking, classical DA algorithms are categorized into filters and smoothers. Filters estimate the state using observations up to and including the present, whereas smoothers incorporate information from the entire observation window, including past, present, and future observations. While filtering is the standard approach for real-time forecasting, smoothing is theoretically more accurate for retrospective analysis (Simon, 2006; Evensen and Van Leeuwen, 2000; Buehner et al., 2017). On the one hand, smoothing systematically improves estimation error and reduces uncertainty (Simon, 2006; Anderson and Moore, 1979). On the other hand, smoothing can mitigate the detection delay that filters often exhibit at the onset of extreme events. This distinction becomes particularly important in systems where the underlying dynamics are driven by hidden forcing mechanisms. As a result, filters tend to detect the driving signal only after its effects have been observed (Chen and Majda, 2016). In contrast, by leveraging future information, smoothers can more accurately identify the onset and persistence of the underlying driving mechanisms (Andreou et al., 2026, 2024; Andreou and Chen, 2025).
@@
Although a wide range of DA algorithms has been developed, many methods face theoretical and practical limitations. The Kalman filter and the Rauch–Tung–Striebel (RTS) smoother provide optimal solutions, but only for linear-Gaussian systems (Kalman, 1960; Rauch et al., 1965). In complex dynamical systems, however, strongly non-Gaussian features such as skewness and heavy tails often violate the Gaussian assumptions underlying the classical theory (Jazwinski, 1970). For general nonlinear dynamics, variational DA approaches typically provide deterministic state estimates that do not quantify uncertainty, while particle methods suffer from severe computational burdens in high dimensions due to the curse of dimensionality (Snyder et al., 2008; Rabier, 2005; Lorenc, 2003). Consequently, ensemble Kalman methods have emerged as an effective alternative (Evensen, 1994; Burgers et al., 1998). By approximating system statistics with a finite ensemble, these methods provide state estimates that explicitly quantify uncertainty while remaining derivative-free and computationally feasible in many high-dimensional settings. In practice, the ensemble size (typically $O(10^{2})$) is often chosen independently of the system’s effective dimension (Evensen, 2003; Kalnay et al., 2007). Due to these advantages, ensemble methods have been widely adopted in large-scale operational applications, including numerical weather prediction, ocean science, geophysics, and fluid dynamics (Houtekamer and Zhang, 2016; Iglesias et al., 2013; Evensen, 2003; Aanonsen et al., 2009). Most existing ensemble filtering and smoothing algorithms, however, are formulated in discrete time, reflecting the sequential structure of forecast-analysis cycles. Efficient continuous-time smoothers for nonlinear systems remain comparatively less explored, despite their importance for systems naturally described by stochastic differential equations. Existing formulations, such as the continuous-time unscented RTS smoother (Särkkä, 2010), directly propagate the smoother mean and covariance. In contrast, our ensemble-based smoother evolves stochastic trajectories, providing conditional samples for downstream tasks such as model discovery, while naturally admitting covariance localization for high-dimensional systems.
@@
In this paper, we develop a continuous-time ensemble Kalman–Bucy smoother (EnKBS) for efficient state estimation and trajectory sampling in complex nonlinear dynamical systems. To address high-dimensional settings, the method incorporates standard regularization techniques, including covariance localization and inflation, which mitigate sampling errors, improve numerical stability, and prevent ensemble divergence in nonlinear models (Anderson, 2001; Bergemann and Reich, 2010a; Houtekamer and Mitchell, 1998). Beyond state estimation, we demonstrate that EnKBS provides a powerful tool for dynamical discovery. We incorporate EnKBS into a Bayesian causal inference framework that identifies causal relationships by tracing driving signals backwards from their observed effects (Andreou et al., 2026). Numerical tests are presented using a nonlinear dyad model exhibiting extreme events, for which closed-form filtering and smoothing distributions are available (Andreou et al., 2026; Andreou and Chen, 2024a). We verify asymptotic convergence of the method to the exact analytical solution and show that it remains accurate and statistically consistent even with an ensemble size of $O(10)$. We further integrate EnKBS into a causality-based iterative learning algorithm for partially observed systems, where EnKBS-based trajectory sampling enables the recovery of hidden parameters and identification of the underlying model structure (Chen and Zhang, 2023).
@@
The rest of this paper is organized as follows. Sect. 1 formulates the problem and introduces the proposed continuous-time EnKBS, together with regularization strategies for high-dimensional systems. Sect. 2 evaluates EnKBS on the high-dimensional Lorenz-96 system (Lorenz, 2006) and presents applications to assimilative causal inference (ACI) and causality-based model discovery (Chen and Zhang, 2023; Andreou et al., 2026). We conclude the paper with a discussion in Sect. 2.3.3. Mathematical derivations of the EnKBS and its connection to the discrete-time RTS smoother in a formal time limit are deferred to Appendix A. Finally, Appendix B shows that the proposed EnKBS recovers the exact mean and covariance equations of the classical continuous-time RTS smoother in the linear-Gaussian setting.
@@ 1.1 Problem Setup
We consider a general class of continuous-time nonlinear systems. Let $\vec{x}(t)\in\mathbb{R}^{n_{\vec{x}}}$ denote the hidden states and $\vec{y}(t)\in\mathbb{R}^{n_{\vec{y}}}$ denote the observations (data). The dynamics are governed by the following stochastic differential equations (SDEs) (Jazwinski, 1970; Särkkä and Solin, 2019):
$\mathrm{d}\vec{x}$ $=\vec{f}(\vec{x},\vec{y},t)\mathrm{d}t+\mathbf{\Sigma}(\vec{y},t)^{1/2}\mathrm{d}\vec{B}(t),$ (1a)
$\mathrm{d}\vec{y}$ $=\vec{h}(\vec{x},\vec{y},t)\mathrm{d}t+\mathbf{\Gamma}(\vec{y},t)^{1/2}\mathrm{d}\vec{W}(t).$ (1b)
$\vec{f}$ is the drift of the signal process Eq. (1a) and $\vec{h}$ is the drift of the observation process Eq. (1b). The terms $\vec{B}(t)\in\mathbb{R}^{n_{\vec{x}}}$ and $\vec{W}(t)\in\mathbb{R}^{n_{\vec{y}}}$ are independent standard Wiener processes. For simplicity, we assume that the signal process noise covariance $\mathbf{\Sigma}(\vec{y},t)\in\mathbb{R}^{n_{\vec{x}}\times n_{\vec{x}}}$ and the observation noise covariance $\mathbf{\Gamma}(\vec{y},t)\in\mathbb{R}^{n_{\vec{y}}\times n_{\vec{y}}}$ are symmetric and positive definite. Throughout the paper, all SDEs, whether written in differential form or in formal derivative form, are understood in the Itô sense (Jazwinski, 1970; Øksendal, 2003).
We consider a fixed time horizon $t\in[0,T]$, and aim to characterize the conditional distribution of $\vec{x}(t)$ given the available observations. Let $\vec{y}_{[0,\tau]}=\{\vec{y}(s):0\leq s\leq\tau\}$ denote the observation trajectory up to time $\tau\in[0,T]$. We formulate the filtering and smoothing problems of DA as follows:
@@
    Filtering: Estimate the conditional distribution of $\vec{x}(t)$ given the observational history up to the present $\vec{y}_{[0,t]}$. The filtering distribution is denoted by $p(\vec{x}(t)\mid\vec{y}_{[0,t]})$, with mean $\overline{\vec{x}}_{\mathrm{f}}(t)$ and covariance $\mathbf{P}_{\mathrm{f}}(t)$.
@@
    Smoothing: Estimate the conditional distribution of $\vec{x}(t)$ given the full observation trajectory $\vec{y}_{[0,T]}$, including the past and the future. The smoothing distribution is denoted by $p(\vec{x}(t)\mid\vec{y}_{[0,T]})$, with mean $\overline{\vec{x}}_{\mathrm{s}}(t)$ and covariance $\mathbf{P}_{\mathrm{s}}(t)$.
@@
It is worth mentioning that the coefficients in Eq. (1), including the drifts $\vec{f}$ and $\vec{h}$, as well as the noise covariances $\mathbf{\Sigma}$ and $\mathbf{\Gamma}$, may depend explicitly on the observation history $\vec{y}_{[0,t]}$. Since $\vec{y}$ is given in the filter and smoother, its trajectory can be treated as time-varying parameters (Crisan and Rozovskii, 2011). This flexibility allows for broad application in which the observed variables actively feed back on and modulate the hidden dynamics; however, for our purposes, we assume the Markovian structure in Eq. (1) (Øksendal, 2003). For notational simplicity, we may omit the explicit dependence of the model components on the dependent and time variables when the context is clear, and write $\vec{f}(\vec{x},\vec{y},t)$ as $\vec{f}$, and similarly for $\vec{h}$, $\mathbf{\Sigma}$, and $\mathbf{\Gamma}$.
@@
In general, evolving the full filtering or smoothing distribution is computationally infeasible (Crisan and Rozovskii, 2011). Therefore, we focus on estimating the first two moments: the mean and covariance. The idea is that in systems with linear dynamics and Gaussian noise, the conditional distributions are Gaussian and fully described by their mean and covariance, i.e., $p(\vec{x}(t)\mid\vec{y}_{[0,t]})=\mathcal{N}(\overline{\vec{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t))$ and $p(\vec{x}(t)\mid\vec{y}_{[0,T]})=\mathcal{N}(\overline{\vec{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t))$. These exact moments are governed by the classical Kalman–Bucy Filter (Kalman and Bucy, 1961) and continuous-time Rauch–Tung–Striebel (RTS) smoother equations (Rauch et al., 1965).
@@
For general nonlinear systems where closed-form DA solutions are unavailable, we adopt an ensemble-based approach (Evensen, 2009, 1994). Here, the conditional distributions are represented by a finite ensemble of size $m$. This can be viewed as a Monte Carlo approximation in which the conditional moments are replaced by empirical ensemble moments. Each ensemble member independently evolves according to Eq. (1), but the ensemble constituents are coupled through the empirical mean and covariance.
@@
Let $\vec{x}^{(i)}(t)$, $i=1,\dots,m$, denote the ensemble members. The empirical mean $\overline{\vec{x}}(t)$ and covariance $\mathbf{P}(t)$ for $t\in[0,T]$ are defined by
$\overline{\vec{x}}(t):=\frac{1}{m}\sum_{i=1}^{m}\vec{x}^{(i)}(t),\qquad\mathbf{P}(t):=\frac{1}{m-1}\sum_{i=1}^{m}\big(\vec{x}^{(i)}(t)-\overline{\vec{x}}(t)\big)\big(\vec{x}^{(i)}(t)-\overline{\vec{x}}(t)\big)^{\top}.$
@@
When needed, we use subscripts $\mathrm{f}$ and $\mathrm{s}$ to distinguish between the filter- and smoother-based ensembles. For brevity, we use $\big(\overline{\vec{x}}_{\mathrm{f}},\mathbf{P}_{\mathrm{f}}\big)$ and $\big(\overline{\vec{x}}_{\mathrm{s}},\mathbf{P}_{\mathrm{s}}\big)$ to denote the corresponding empirical moments, and do not distinguish them from the true moments unless confusion may arise.
@@
While filtering is the standard approach for online state estimation, many applications, such as Earth system reanalysis (Kalnay, 2003) and parameter estimation (Navon, 1998; Yang and Delsole, 2009), necessitate a retrospective analysis once the full observation trajectory over $[0,T]$, $\vec{y}_{[0,T]}$, is available. Based on the filtering quantities $\big(\overline{\vec{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t)\big)$, the smoothing quantities $\big(\overline{\vec{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t)\big)$ incorporate information from future observations $\vec{y}_{[t,T]}$, producing more accurate estimates and reducing uncertainty over $[0,T]$. This improvement motivates a systematic forward-backward structure algorithm (Evensen, 2009; Särkkä and Svensson, 2023):
@@
    Forward pass (filtering): Integrate forward from $t=0$ to $t=T$ to compute the filter ensemble $\vec{x}^{(i)}_{\mathrm{f}}(t),i=1,\dots,m$ (forecasting). The filtering mean and covariance $\{\overline{\vec{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t)\}_{0\leq t\leq T}$ are obtained.
@@
    Backward pass (smoothing): Integrate backward-in-time from $t=T$ to $t=0$. Using the filtering estimates as a prior, this pass produces the smoothing ensemble $\vec{x}_{\mathrm{s}}^{(i)}(t),i=1,\dots,m$ and therefore its statistics $\{\overline{\vec{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t)\}_{0\leq t\leq T}$ (hindcasting).
In our method, we implement this architecture by utilizing the continuous-time ensemble Kalman–Bucy filter (EnKBF) for the forward pass (Bergemann and Reich, 2012), and the EnKBS for the backward pass. The operational details of these methods are presented in the following subsections.